News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Monday 22 May 2000 8.30am EDT

During the last week, Chandra completed the observing schedule as planned.

This week's loads were rebuilt on 5/15 after noting a failure to acquire 3 of the 5 acquisition stars for an observation (Q0059-2735) on 5/15. Acquisition Stars had also been dropped by the Aspect Camera on a number of other observations over the weekend. If the number of stars acquired falls below 2 then the spacecraft will stop the daily load and transition to bright star hold. Engineering assessment indicated that the aspect camera performance had not changed, and that the likely cause was a recent change to the parameters used by the OFLS in determining the faint limit for acquisition stars. The change was made in conjunction with the installation of a new star catalog and was based on analysis of prior on-orbit data. The loads for the remaining part of the week were rebuilt using an increased error term to avoid the possibility of a bright star hold. The loads were uplinked without impact to spacecraft operations, and the remaining acquisitions for the week were made as expected. Further analysis is underway to determine the optimal error terms for use by the OFLS.

The first HETG insertion and retraction commanded from the daily load using the B-side MCE was completed successfully for the 5/18 observation of Ex Hya. Telemetry indicated that both the insertion and retraction were nominal. Operations will continue using the B-side MCE while other approaches for returning A-side redundancy are assessed.

Preparations were completed for the 4th ACIS Squeegee mode test scheduled for 5.30pm EDT on 5/22. This test will include initial flushes of the CCD to remove charge built up while the CCD is not being clocked in order to remove a bias instability seen during the earlier on-orbit tests. Two runs will be performed: a 16 row squeegee with a 32 row exclude window, and a control run.

The planned observations for the coming week are as follows.

------------------------------------------
PSS 1435+3057          ACIS-S          May 21
Cen Cluster            ACIS-S
BHXB (Outburst Decay)  ACIS-S/HETG     May 22
Radiation Belts
NGC 6652               HRC-I
NGC 6441               HRC-I
Ophiuchi Mosaic #4     ACIS-I          May 23
1735-269               ACIS-S
G21.5-0.9 (a series)   ACIS-S
NGC 5846               ACIS-S          May 24
Radiation Belts
Abell 2104             ACIS-S          May 25
1217+301               HRC-I           May 26
Cygnus A               ACIS-S
M31 (N, S, Center)     HRC-I
Q2345+007              ACIS-S
AXAF Southern Deep Fld ACIS-I
Radiation Belts                        May 27
E0102-72.3             ACIS-S          May 28
Q1127-145              ACIS-S

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports